OpenStudy (anonymous):

Jose needs to enclose a rectangular section of his yard. The perimeter of the section is 27 feet, and the area is 35 square feet. Find the length and width of the section. Part I: Let L = length of the section and let W = width of the section. Write an equation for the perimeter of the section. Hint: P = 2(L+W). Part II: Use the equation you wrote in Step 1 and solve for W. Show your work. Part III: Using L and W as defined in Step 1, write an equation for the area of the section. Hint: A = LW. Part IV: Substitute the expression you found for W in Part 2 into the area equation in Part 3.

OpenStudy (lncognlto):

Okay, do you know how to begin?

OpenStudy (anonymous):

not really

OpenStudy (lncognlto):

Right, p = 2(l + w), as given. so 27 = 2(l+w). Now isolate w.

OpenStudy (anonymous):

how would i do that?

OpenStudy (lncognlto):

Just get it on its own: 27/2 = l + w so w = ?

OpenStudy (anonymous):

13.5?

OpenStudy (lncognlto):

13.5 - l, exactly. Then A = LW so 35 = lw

OpenStudy (anonymous):

thats just the area 35

OpenStudy (lncognlto):

Yes, now substituting w = 13.5 - 1 into lw = 35 gives l(13.5-l) = 35 sorry, I got lost.

OpenStudy (lncognlto):

Sorry, I can't make this work. You can repost the question to get some proper help :)

OpenStudy (anonymous):

okay